OpenStudy (anonymous):

Find f(g(x)). f(x)= sqrt of x+3 g(x) = 8x - 7

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

is f(x) the function \[\Large f(x) = \sqrt{x}+3\] or is it \[\Large f(x) = \sqrt{x+3}\]

OpenStudy (anonymous):

2nd one

jimthompson5910 (jim_thompson5910):

first replace every x with g(x) \[\Large f(x) = \sqrt{x+3}\] \[\Large f(g(x)) = \sqrt{g(x)+3}\]

jimthompson5910 (jim_thompson5910):

then on the right side, replace g(x) with 8x-7 (since g(x) = 8x-7) and simplify

jimthompson5910 (jim_thompson5910):

so our full step by step work looks like this \[\Large f(x) = \sqrt{x+3}\] \[\Large f(g(x)) = \sqrt{g(x)+3}\] \[\Large f(g(x)) = \sqrt{8x-7+3}\] \[\Large f(g(x)) = \sqrt{8x-4}\]

jimthompson5910 (jim_thompson5910):

I guess you could pull out a 4 and simplify, but I think this is far enough for the teacher to be satisfied

OpenStudy (anonymous):

I think it is supposed to be simplified So if i simplify it, it would be 2 sqrt of 2x-1?

jimthompson5910 (jim_thompson5910):

Good, \[\Large f(g(x)) = \sqrt{8x-4}\] \[\Large f(g(x)) = \sqrt{4(2x-1)}\] \[\Large f(g(x)) = \sqrt{4}*\sqrt{2x-1}\] \[\Large f(g(x)) = 2\sqrt{2x-1}\]

OpenStudy (anonymous):

Thank you!! I understand it now

jimthompson5910 (jim_thompson5910):

I'm glad you do

OpenStudy (anonymous):

Can you help me with one more problem i'm stuck on?

jimthompson5910 (jim_thompson5910):

sure